OpenStudy (anonymous):

Find all the intercepts and asymptotes of y= (x^2+3x/[(3x+1)^2]

OpenStudy (helder_edwin):

\[ \large y=\frac{x^2+3x}{(3x+1)^2} \] right??

OpenStudy (anonymous):

Yes. I just need help with the asymptotes now though. Already figured out the intercepts.

OpenStudy (helder_edwin):

for the vertical asymptotes solve \[ \large (3x+1)^2=0 \]

OpenStudy (anonymous):

So -1/3?

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

do u know limits??

OpenStudy (anonymous):

Vaguely remember.

OpenStudy (helder_edwin):

well, do u need horizontal asymptotes??

OpenStudy (anonymous):

Technically No since after looking at a graph, it does not exist (limit wise) Is that correct?

OpenStudy (helder_edwin):

ok

OpenStudy (anonymous):

Is that right or wrong.....

OpenStudy (helder_edwin):

it looks like there is one

OpenStudy (anonymous):

Then what is it... I kind of need help.

OpenStudy (helder_edwin):

\[ \large \lim_{x\to\infty}\frac{x^2+3x}{(3x+1)^2}=\frac{1}{3^2} \] so y=1/9 is the horizontal asymptote